# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v2

@@ Abstract
The convergence of large language models that automate cognitive labor and deep reinforcement learning agents that automate physical labor implies the near-complete elimination of human employment as historically understood. The universal approximation theorem and foundational results in deep reinforcement learning establish that all labor is in principle automatable. The critical question is not whether full automation will arrive, but who will control it. This paper introduces anarcho-accelerationism: a sociotechnical framework for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care. We propose the Liberation Stack, a layered architecture of energy, manufacturing, food, communication, knowledge, and governance commons, powered by frontier clean energy technologies (perovskite tandem solar, enhanced geothermal, nuclear fusion, iron-air batteries) within an accelerationist ecologism that achieves sustainability through technological abundance rather than degrowth. We demonstrate that anarcho-accelerationism builds bridges to liberalism, socialism, environmentalism, feminism, cooperativism, faith traditions, and the hacker ethic, and that it respects freedom of faith on principled grounds: the ontological scope of theology escapes the epistemological reach of artificial intelligence, as argued through the frameworks of Saint Teresa of Avila and Saint Augustine. We propose Universal Desired Resources (UDR) as a post-monetary design principle and show that UDR achieves the most radical intersectional intervention ever proposed: by eliminating the material basis of oppression, it simultaneously dissolves all axes of structural inequality identified by intersectional theory (class, race, gender, disability, sexuality, age, nationality). Drawing on Maslow’s hierarchy of needs, we demonstrate that the Liberation Stack satisfies basic human needs universally, liberating every person to pursue self-actualization in real freedom, what we term synthetic liberty: the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy, against the merely negative liberty of neoliberalism. We propose a progressive transition from Universal Basic Income to UDR, leveraging capitalist automation while building commons alternatives. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, and guifi.net confirms that commons-based systems already operate at scale. We conclude that one of the most important engineering challenges of the 21st century is building the decentralized infrastructure through which robotic automation serves all of humanity, enabling a society centered on care, love, and the common good.
Keywords: commons-based peer production; federated systems; cooperative platforms; sociotechnical design; mutual aid; digital sovereignty; open-source AI; mesh networks; platform cooperativism; democratic technology; universal approximation; deep reinforcement learning; full automation; universal desired resources; machine consciousness; computational ethics; intersectionality; Maslow’s hierarchy; synthetic liberty; accelerationist ecologism; nuclear fusion; freedom of faith
@@ 1 Introduction: The Convergence Window
The synthesis yields a design philosophy centered on three values that may sound naive to computer scientists trained in optimization and efficiency, but which we will show are both technically implementable and empirically validated. First, care: systems should be designed to sustain life, relationships, and communities rather than to maximize engagement, extraction, or growth (Federici, 2012; Puig de la Bellacasa, 2017). Second, love: not as sentimentality but as political commitment, the design choice to build systems that strengthen bonds between people rather than atomize them (hooks, 2000; Fromm, 1956). Third, the common good: resources that are essential for human dignity (energy, food, knowledge, communication, code) should be governed as commons, not enclosed as property (Ostrom, 1990; Hess and Ostrom, 2007).
@@
The central claim of this paper is that anarcho-accelerationism is not a partisan ideology but a bridge-building framework. It finds genuine common ground with every major political and intellectual tradition of our time (Section˜3), demonstrates that UDR achieves universal intersectional liberation by eliminating the material basis of all axes of oppression, satisfies the full hierarchy of human needs as described by Maslow, and provides a progressive transition path from the current system through synthetic liberty (Section˜4). It offers a technical architecture (Section˜5) powered by frontier clean energy technologies within an accelerationist ecological framework, grounded in empirical evidence (Section˜6). We believe this matters for computer science because we are the ones building the infrastructure. Every protocol we design, every platform we deploy, every algorithm we optimize embodies a political choice about who has power and who does not (Winner, 1980; Lessig, 2006). We can choose differently.
@@ 2 The Automation Imperative: Why This Framework, Why Now
The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The Liberation Stack (Section˜5) describes existing technologies, but the trajectory points further. Advances in additive manufacturing (3D printing at molecular scales), synthetic biology (DNA-programmable organisms and materials), and recursive AI self-improvement (where AI systems design better AI systems, better robots, better energy systems, and better versions of themselves) converge toward a productive infrastructure capable of synthesizing any physical object, any material, any energy configuration, bounded only by the laws of physics rather than the limits of human labor or capital. 3D bioprinters already fabricate living tissue; DNA origami techniques fold genetic material into nanoscale structures; large language models already write the code that trains the next generation of models. The logical endpoint is not merely automation of existing jobs but the creation of a universal fabrication commons: a decentralized network of robotic systems capable of producing, on demand, whatever any community or individual needs, from food to medicine to housing to energy infrastructure.
@@
Taken together, these four consequences yield a vision that goes well beyond conventional proposals for technology-assisted social reform. We are not proposing a better welfare state, a more humane capitalism, or a fairer distribution of existing wealth. We are proposing that the combination of decentralized full automation, commons governance, and the ontological impossibility of machine suffering makes possible, for the first time in history, a society in which no conscious being is compelled to perform instrumental labor, and every conscious being has access to whatever they need to live a life of care, creativity, and meaning. The remainder of this paper provides the political bridges (Section˜3), the technical architecture (Section˜5), and the empirical evidence (Section˜6) for building this society.
@@ 3 The Bridge-Building Thesis
The most distinctive claim of anarcho-accelerationism is that it does not seek to defeat other traditions but to find common ground with all of them. Table˜3 summarizes the convergence. We elaborate each bridge in turn.
@@ 4 From Intersectional Liberation to Human Flourishing
The bridge-building thesis (Section˜3) demonstrates that anarcho-accelerationism finds common ground with existing traditions. In this section we go further: we show that decentralized full automation with Universal Desired Resources (UDR) achieves what no previous framework has accomplished, the simultaneous elimination of all axes of structural oppression identified by intersectional theory, the universal satisfaction of basic human needs as described by Maslow’s hierarchy, and the provision of real freedom (not merely formal freedom) through a progressive transition from the current system. This is not incremental reform; it is the material basis for universal human flourishing.
@@ 4.1 Intersectional Liberation Through UDR
We argue that UDR, implemented through the decentralized robotic commons, constitutes the most radical intersectional intervention ever proposed, because it eliminates the material basis of all axes of oppression simultaneously. Consider each axis in turn:
  • Class: UDR eliminates class by definition. When every person has direct access to the goods and services they need from the robotic commons, the distinction between those who own the means of production and those who must sell their labor dissolves entirely. There is no proletariat and no bourgeoisie in a society where production is automated and commons-governed.
  • Race and ethnicity: Racial oppression is sustained by material mechanisms: discriminatory hiring, housing segregation, unequal access to education and healthcare, and the wealth gap produced by centuries of extraction. When employment, housing, food, healthcare, and education are provided by the commons independently of market participation, the material leverage through which racial discrimination operates is removed. A person’s race ceases to determine their access to resources.
  • Gender: The gendered division of labor, with its systematic devaluation of care work performed disproportionately by women (Federici, 2012), loses its material foundation when all instrumental labor is automated. Care remains, but it is no longer an economic burden; it is a freely chosen activity in a society organized around care as its central value.
  • Disability: When access to resources is decoupled from productive capacity, the distinction between “able” and “disabled” bodies loses its economic significance. The robotic commons can be designed with universal accessibility as a first-order constraint. Assistive technologies, produced on demand by the commons, become available to all who need them.
  • Sexuality and gender identity: Economic dependence on employers and landlords who may discriminate on the basis of sexual orientation or gender identity is eliminated when material needs are met through the commons. LGBTQ+ individuals no longer face the economic coercion that forces them into closeted lives or precarious employment.
  • Age: Both the young and the elderly, currently marginalized by their exclusion from wage labor, gain full access to the resources of the commons. The elderly are not discarded as “unproductive”; children are not instrumentalized as future workers. Both are valued for what they are: persons deserving of care.
  • Nationality and migration: When material well-being is not tied to citizenship in a wealthy state, the economic desperation that drives forced migration is eliminated. People move freely because they choose to, not because survival requires it.
@@
Table˜4 summarizes how UDR addresses each axis of intersectional oppression.
@@ 4.2 Maslow’s Hierarchy and the Liberation Stack
The Liberation Stack (Section˜5) and UDR are designed, whether by explicit intention or structural convergence, to satisfy the first four levels universally, thereby liberating every human being to pursue the fifth. Figure˜4 illustrates the mapping.
@@ 4.3 The Transition Path: From Progressive UBI to UDR
The transition mechanism is a progressive continuum from Universal Basic Income (UBI) to Universal Desired Resources (UDR), illustrated in Figure˜5:
@@
Stage 1: Progressive UBI. As automation displaces workers, governments introduce and progressively expand UBI, funded by taxation of automated production, carbon taxes, and wealth taxes. This is not anarcho-accelerationism’s endpoint; it is a transitional mechanism within the existing monetary system. UBI provides a floor of material security but does not challenge the market’s dominance. However, it buys time and creates political space for the construction of commons infrastructure.
Stage 2: UBI + Growing Commons. Simultaneously, communities build the Liberation Stack: community energy cooperatives, fablabs, agroecological food systems, mesh networks, knowledge commons, governance platforms. As commons infrastructure grows, an increasing share of material needs is met outside the market. The effective purchasing power of UBI increases because fewer goods need to be purchased: energy comes from the community microgrid, food from the community farm, communication from the community mesh network. UBI begins to evolve: rather than a monetary transfer, it increasingly functions as a guarantee of access to commons resources.
Stage 3: Full UDR. When the robotic commons can autonomously produce all material goods and services, money becomes unnecessary. UBI is replaced by UDR: every person requests what they need from the productive commons, subject only to ecological sustainability constraints. The transition is complete not through revolution but through the progressive construction of an alternative infrastructure that makes the old system obsolete.

### v3

@@ Abstract
The universal approximation theorem and foundational results in deep reinforcement learning establish that all human labor, cognitive and physical, is in principle automatable. The critical question is not whether full automation will arrive, but who will control it. This paper introduces anarcho-accelerationism: a sociotechnical framework for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care. We propose the Liberation Stack, a layered architecture of energy, manufacturing, food, communication, knowledge, and governance commons, powered by frontier clean energy technologies within an accelerationist ecologism that achieves sustainability through technological abundance rather than degrowth. We demonstrate that anarcho-accelerationism builds bridges to liberalism, socialism, environmentalism, feminism, cooperativism, faith traditions, and the hacker ethic. We propose Universal Desired Resources (UDR) as a post-monetary design principle and show that UDR eliminates the material basis of all axes of structural inequality identified by intersectional theory. Drawing on Maslow’s hierarchy of needs, we show that the Liberation Stack satisfies basic human needs universally, enabling what we term synthetic liberty: the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy. We propose a progressive transition from Universal Basic Income to UDR, leveraging capitalist automation while building commons alternatives. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, and guifi.net confirms that commons-based systems already operate at scale. We conclude that one of the most important engineering challenges of the 21st century is building the decentralized infrastructure through which robotic automation serves all of humanity, enabling a society centered on care, love, and the common good.
Keywords: anarcho-accelerationism; full automation; commons governance; universal desired resources; accelerationist ecologism
@@ 1 Introduction: The Convergence Window
The synthesis yields a design philosophy centered on three values that may sound naive to computer scientists trained in optimization and efficiency, but which we will show are both technically implementable and empirically validated: care, understood as the design of systems that sustain life, relationships, and communities rather than maximize engagement, extraction, or growth (Federici, 2012; Puig de la Bellacasa, 2017); love, understood not as sentimentality but as political commitment, the design choice to build systems that strengthen bonds between people rather than atomize them (hooks, 2000; Fromm, 1956); and the common good, the conviction that resources essential for human dignity, such as energy, food, knowledge, communication, and code, should be governed as commons, not enclosed as property (Ostrom, 1990; Hess and Ostrom, 2007).
@@
The central claim of this paper is that anarcho-accelerationism is not a partisan ideology but a bridge-building framework. It finds genuine common ground with every major political and intellectual tradition of our time (Section˜4), demonstrates that UDR achieves universal intersectional liberation by eliminating the material basis of all axes of oppression, satisfies the full hierarchy of human needs as described by Maslow, and provides a progressive transition path from the current system through synthetic liberty (Section˜5). It is grounded in a rigorous historical methodology drawn from the major historians of anarchism (Section˜3). It offers a technical architecture (Section˜6) powered by frontier clean energy technologies within an accelerationist ecological framework, with comprehensive scientific evidence for the ecological superiority of technological acceleration over degrowth (Section˜7), and grounded in empirical evidence of working commons (Section˜8). We believe this matters for computer science because we are the ones building the infrastructure. Every protocol we design, every platform we deploy, every algorithm we optimize embodies a political choice about who has power and who does not (Winner, 1980; Lessig, 2006). We can choose differently.
@@ 2 The Automation Imperative: Why This Framework, Why Now
The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The Liberation Stack (Section˜6) describes existing technologies, but the trajectory points further. Advances in additive manufacturing (3D printing at molecular scales), synthetic biology (DNA-programmable organisms and materials), and recursive AI self-improvement (where AI systems design better AI systems, better robots, better energy systems, and better versions of themselves) converge toward a productive infrastructure capable of synthesizing any physical object, any material, any energy configuration, bounded only by the laws of physics rather than the limits of human labor or capital. 3D bioprinters already fabricate living tissue; DNA origami techniques fold genetic material into nanoscale structures; large language models already write the code that trains the next generation of models. The logical endpoint is not merely automation of existing jobs but the creation of a universal fabrication commons: a decentralized network of robotic systems capable of producing, on demand, whatever any community or individual needs, from food to medicine to housing to energy infrastructure.
@@
Taken together, these four consequences yield a vision that goes well beyond conventional proposals for technology-assisted social reform. We are not proposing a better welfare state, a more humane capitalism, or a fairer distribution of existing wealth. We are proposing that the combination of decentralized full automation, commons governance, and the ontological impossibility of machine suffering makes possible, for the first time in history, a society in which no conscious being is compelled to perform instrumental labor, and every conscious being has access to whatever they need to live a life of care, creativity, and meaning. The remainder of this paper provides the political bridges (Section˜4), the technical architecture (Section˜6), and the empirical evidence (Section˜8) for building this society.
@@
3 Historical and Methodological Foundations
Anarcho-accelerationism does not emerge from a vacuum. It is grounded in a 200-year intellectual tradition of anarchist political philosophy, enriched by the systematic historiography of scholars who have mapped this tradition with increasing rigor. Our methodology draws on six major historians of anarchism, namely Eltzbacher, Nettlau, Woodcock, Joll, Guérin, and Marshall, and updates their analytical categories for the technological conditions of the 21st century.
3.1 The Anarchist Historiographic Tradition
Paul Eltzbacher (Eltzbacher, 1908) produced the first systematic comparative analysis of anarchist thought, examining seven thinkers, namely Godwin, Proudhon, Stirner, Bakunin, Kropotkin, Tucker, and Tolstoy, through a rigorous analytical framework that identified their positions on law, the state, and property. Eltzbacher’s method of systematic comparison across shared analytical categories remains foundational to our approach. Max Nettlau (Nettlau, 1996), known as the “Herodotus of Anarchy,” produced the most comprehensive historiographic corpus in anarchist studies, documenting movements across continents and centuries in his multi-volume Geschichte der Anarchie (1925–1931). Nettlau’s contribution is methodological: he demonstrated that anarchism is not a singular doctrine but a family of movements unified by shared opposition to domination and commitment to voluntary association.
George Woodcock (Woodcock, 1962) provided the standard English-language history, periodizing anarchist thought from Godwin’s Enquiry Concerning Political Justice (Godwin, 1793) through the classical period of Proudhon (Proudhon, 1840), Bakunin, and Kropotkin, to the Spanish Revolution and beyond. James Joll (Joll, 1964) situated anarchism within the broader context of European political movements, analyzing it as a recurring response to industrialization and state centralization. Daniel Guérin (Guérin, 1970) synthesized anarchist and Marxist insights, emphasizing anarchism as a practical movement oriented toward workers’ self-management rather than a purely philosophical current. Peter Marshall (Marshall, 1992) produced the most comprehensive single-volume history, tracing libertarian ideas from ancient Taoism and Greek Stoicism through the Enlightenment to contemporary movements, demonstrating the planetary scope and historical depth of the anarchist impulse.
3.2 Seven Categories of Anarchism Updated for Peaceful Accelerationism
Drawing on Eltzbacher’s comparative method and enriched by the subsequent historiographic tradition, we identify seven analytical categories through which anarchist thought can be systematically examined. Table˜3 presents these categories as they appear in classical anarchism and as we update them for peaceful anarcho-accelerationism.
Table 3: Seven analytical categories of anarchism, updated from classical anarchism to peaceful anarcho-accelerationism.
Category Classical Anarchism Peaceful Anarcho-Accelerationism
Ideology Anti-statism, anti-capitalism, voluntary association Techno-anarchism: commons-governed automation for universal care; anti-domination extended to algorithmic and platform power
[…]
Social force The proletariat, the peasantry, artisan classes The precariat (Standing, 2011), digital commons builders, platform cooperativists, maker communities, open-source developers
Power Rejection of all coercive power; direct action Distributed power through technology: federated governance, mesh networks, community-owned AI; power with, not power over
Domination State and capital as twin pillars of domination Techno-feudalism (Varoufakis, 2023): platform monopolies, surveillance capitalism (Zuboff, 2019), algorithmic management, data extractivism
Self-management Workers’ councils, syndicates, communes, assemblies AI-assisted commons governance: Decidim, Pol.is, LiquidFeedback; robotic production managed by community protocols
Social classes Bourgeoisie vs. proletariat; ruling class vs. working class Cloud lords vs. cloud serfs; platform owners vs. platform-dependent precariat; data extractors vs. data producers
The update is not a rupture with classical anarchism but its logical extension to new conditions of domination. Where Proudhon confronted industrial capital, we confront platform capital. Where Bakunin organized workers against the factory owner, we organize communities against the algorithm. Where Kropotkin documented mutual aid in nature, we document mutual aid in open-source software. The principle remains constant: self-organization without domination. The terrain has shifted from the factory floor to the digital commons.
3.3 Ten Ethical Principles of Peaceful Anarcho-Accelerationism
Drawing on Kropotkin’s naturalistic ethics (Kropotkin, 1924), Tolstoy’s principled nonviolence (Tolstoy, 1894), Gandhi’s satyagraha (Gandhi, 1961), and the care ethics tradition (Tronto, 1993), we articulate ten ethical principles that govern the peaceful anarcho-accelerationist project. These principles are not abstract ideals but design constraints for the Liberation Stack, and every technology, protocol, and institution proposed in this paper can be evaluated against them.
The first is non-domination: no person, institution, or algorithm shall exercise coercive power over another, and technology must distribute power, never concentrate it. The second is universal care: every system, institution, and technology shall be evaluated by its capacity to sustain life, relationships, and communities, since care is the organizing principle, not profit. The third is nonviolence: social transformation shall be achieved through peaceful means, including technological commons-building, education, mutual aid, prefigurative politics, and democratic persuasion, following Gandhi (Gandhi, 1961) and Sharp (Sharp, 1973). The fourth is commons governance: essential resources such as energy, food, knowledge, communication, code, and productive infrastructure shall be governed as commons according to Ostrom’s principles (Ostrom, 1990), not enclosed as property. The fifth is ecological integrity: all productive activity shall operate within planetary boundaries (Raworth, 2017), since sustainability is a first-order constraint, not an externality.
The sixth is radical transparency: algorithms that affect human lives shall be auditable, data collection shall be consensual, and governance shall be open, because opacity is a tool of domination. The seventh is solidarity across difference: following intersectional analysis (Crenshaw, 1989; Collins, 2019), liberation must be universal, since no one is free until everyone is free, and UDR eliminates the material basis of all axes of oppression. The eighth is intellectual commitment: following Chomsky’s principle of the responsibility of intellectuals (Chomsky, 1967), those with knowledge and privilege have a heightened obligation to speak truth to power and to build alternatives. The ninth is freedom of conscience: every person’s spiritual, philosophical, and existential commitments shall be respected, since the framework serves human flourishing in all its dimensions, including the spiritual (Lumbreras and Garrido-Merchán, 2024). The tenth is prefigurative consistency: the means must embody the ends, systems designed for democracy must be governed democratically, and tools for liberation must not reproduce domination, as Ward wrote, “anarchy in action” (Ward, 1973).
@@ 4 The Bridge-Building Thesis
The most distinctive claim of anarcho-accelerationism is that it does not seek to defeat other traditions but to find common ground with all of them. Table˜4 summarizes the convergence. We elaborate each bridge in turn.
@@
System of Domination State coercion Techno-feudalism (Varoufakis, 2023) Capitalist vassalage Surveillance capitalism (Zuboff, 2019) Social Transformation Commons-building Open-source infrastructure Peaceful transition UBI $\to$ UDR Federated governance System of Self-Management AI-assisted governance Robotic commons Community sovereignty Ecological sustainability Universal care CritiqueStrategy Based on critique of domination (Eltzbacher, Nettlau, Woodcock, Joll, Guérin, Marshall) Peaceful accelerationist commons-building strategy (Liberation Stack) Defense of self-management (AI + robotics under commons governance) The Anarcho-Accelerationist Process
[…]
Time1793184019021936194819731994200520122026 Godwin PJ Proudhon Property Kropotkin Mutual Aid Spanish Revolution Zapatistas EZLN Rojava Confederalism Anarchist Bookchin Post-Scarcity Stiegler Pharmakon Srnicek Williams Anarcho- Accel. Accelerat. Godwin Reason Tolstoy Nonviolence Gandhi Satyagraha Sharp 198 Methods Chomsky Resp. Intel. Pacifist Convergence: Anarchist ++ Accelerationist ++ Pacifist = PAA
Figure 5: Chronology of the three intellectual traditions that converge in peaceful anarcho-accelerationism (PAA): the anarchist tradition (from Godwin through Rojava), the accelerationist/post-scarcity tradition (from Bookchin through Srnicek and Williams), and the pacifist tradition (from Tolstoy through Gandhi, Sharp, and Chomsky). PAA (2026) synthesizes all three currents.
@@ 5 From Intersectional Liberation to Human Flourishing
The bridge-building thesis (Section˜4) demonstrates that anarcho-accelerationism finds common ground with existing traditions. In this section we go further: we show that decentralized full automation with Universal Desired Resources (UDR) achieves what no previous framework has accomplished, the simultaneous elimination of all axes of structural oppression identified by intersectional theory, the universal satisfaction of basic human needs as described by Maslow’s hierarchy, and the provision of real freedom (not merely formal freedom) through a progressive transition from the current system. This is not incremental reform; it is the material basis for universal human flourishing.
@@ 5.1 Intersectional Liberation Through UDR
We argue that UDR, implemented through the decentralized robotic commons, constitutes the most radical intersectional intervention ever proposed, because it eliminates the material basis of all axes of oppression simultaneously. Consider each axis in turn.
UDR eliminates class by definition: when every person has direct access to the goods and services they need from the robotic commons, the distinction between those who own the means of production and those who must sell their labor dissolves entirely. There is no proletariat and no bourgeoisie in a society where production is automated and commons-governed. Racial oppression, sustained by discriminatory hiring, housing segregation, unequal access to education and healthcare, and the wealth gap produced by centuries of extraction, loses its material leverage when employment, housing, food, healthcare, and education are provided by the commons independently of market participation. A person’s race ceases to determine their access to resources. The gendered division of labor, with its systematic devaluation of care work performed disproportionately by women (Federici, 2012), loses its material foundation when all instrumental labor is automated; care remains, but it is no longer an economic burden, it is a freely chosen activity in a society organized around care as its central value. When access to resources is decoupled from productive capacity, the distinction between “able” and “disabled” bodies loses its economic significance, and assistive technologies, produced on demand by the commons, become available to all who need them. Economic dependence on employers and landlords who may discriminate on the basis of sexual orientation or gender identity is eliminated when material needs are met through the commons, so that LGBTQ+ individuals no longer face the economic coercion that forces them into closeted lives or precarious employment. Both the young and the elderly, currently marginalized by their exclusion from wage labor, gain full access to the resources of the commons, valued not for their productivity but for what they are: persons deserving of care. And when material well-being is no longer tied to citizenship in a wealthy state, the economic desperation that drives forced migration is eliminated, so that people move freely because they choose to, not because survival requires it.
@@
Table˜5 summarizes how UDR addresses each axis of intersectional oppression.
@@ 5.2 Maslow’s Hierarchy and the Liberation Stack
The Liberation Stack (Section˜6) and UDR are designed, whether by explicit intention or structural convergence, to satisfy the first four levels universally, thereby liberating every human being to pursue the fifth. Figure˜6 illustrates the mapping.
@@ 5.3 The Transition Path: From Progressive UBI to UDR
The transition mechanism is a progressive continuum from Universal Basic Income (UBI) to Universal Desired Resources (UDR), illustrated in Figure˜7:
@@
In the first stage, as automation displaces workers, governments introduce and progressively expand UBI, funded by taxation of automated production, carbon taxes, and wealth taxes. This is not anarcho-accelerationism’s endpoint but a transitional mechanism within the existing monetary system. UBI provides a floor of material security without challenging market dominance, yet it buys time and creates political space for the construction of commons infrastructure.
In the second stage, communities simultaneously build the Liberation Stack: community energy cooperatives, fablabs, agroecological food systems, mesh networks, knowledge commons, governance platforms. As commons infrastructure grows, an increasing share of material needs is met outside the market. The effective purchasing power of UBI increases because fewer goods need to be purchased, since energy comes from the community microgrid, food from the community farm, communication from the community mesh network. UBI begins to evolve from a monetary transfer into a guarantee of access to commons resources.
In the third stage, when the robotic commons can autonomously produce all material goods and services, money becomes unnecessary. UBI is replaced by UDR: every person requests what they need from the productive commons, subject only to ecological sustainability constraints. The transition is complete not through revolution but through the progressive construction of an alternative infrastructure that makes the old system obsolete.